# Surrogate Modeling for Scalable Evaluation of Distributed Computing Systems for HEP Applications

### v1

@@ 1 Introduction
High energy physics at the LHC relies on the global-scale federated computing infrastructure provided by the Worldwide LHC Computing Grid (WLCG) for processing and storing the sheer amounts of scientific data gathered by the LHC experiments. To achieve the high-throughput demands with limited financial resources, good performance of the computing system has to be constantly ensured. With increasing demands expected from future LHC operations and technical and strategic changes in the federation, the best design for the WLCG infrastructure still needs to be identified.
@@
However, there is a general trade-off between accuracy and the computational complexity, and therefore execution time, that comes with designing practical simulation models as simulations with higher accuracy generally implement more fine-granular simulation models with greater detail. This typically leads to superlinear scaling in the execution time of the respective simulation tools with respect to the size of the simulated infrastructure. As a result, it quickly becomes unfeasible to simulate workflow executions on large infrastructures.
@@ 2 Data Generator DCSim
It has been shown that after calibration, DCSim can achieve valid descriptions of real-world computing workflow executions and systems Horzela2023Proceeding ; Horzela2023_1000165566 . However, the execution time of DCSim scales super-linear with the size of the simulated computing system, which limits its feasibility and necessitates workarounds in studies concerning computing systems at a global scale, like the WLCG. We therefore utilize DCSim as a baseline for testing ML surrogates in simulation.
@@ 3.2 Model Training
We split the data into training and evaluation datasets to train the model, using a 70:30 ratio per simulation of the same length. We employ sequence-to-sequence prediction, predicting exactly one output per input row. The MSE loss function is used to penalize larger errors more heavily than smaller ones, emphasizing the importance of minimizing significant discrepancies between predicted and actual values.
@@ 4.1 Homogeneous Jobs
All models exhibit a negative R-squared value for compute_time, indicating a worse fit than a naive baseline. The KDE plots show that the target distribution exhibits two prominent spikes for the compute_time. While the BiGRU and BiLSTM models successfully capture the general structure of the distribution, they fail to emphasize these two spikes sufficiently. The transformer model, however, predicts mainly the mean values for compute time, failing to capture the actual distribution. While the R-squared value for input_files_transfer_time is positive, it is close to zero for all models, indicating only a slightly better fit than the naive baseline. The KDE plots show that the models capture some distribution features again but fail to emphasize the actual form. In summary, all models failed to predict the observables beyond the first statistical moment, potentially due to the lack of input features providing additional context constraining the job execution, particularly information about the platform architecture. Although the size of the modelled infrastructure is rather minimal in this example, the surrogate ansatz already leads to improvements in the execution time of minimum two orders of magnitude on our test systems ($\mathcal{O}(10\,\mathrm{s})\to\mathcal{O}(100\,\mathrm{ms})$).

### v2

@@ 1 Introduction
High energy physics at the LHC relies on the global-scale federated computing infrastructure provided by the Worldwide LHC Computing Grid (WLCG) wlcg-web for processing and storing the sheer amounts of scientific data gathered by the LHC experiments. To achieve the high-throughput demands with limited financial resources, good performance of the computing system has to be constantly ensured. With increasing demands expected from future LHC operations and technical and strategic changes in the federation, the best design for the WLCG infrastructure still needs to be identified.
@@
However, there is a general trade-off between accuracy and the computational complexity, and therefore execution time, that comes with designing practical simulation models as simulations with higher accuracy generally implement more fine-granular models with greater detail. This typically leads to superlinear scaling in the execution time of the respective simulation tools with respect to the size of the simulated infrastructure. As a result, it quickly becomes unfeasible to simulate workflow executions on large infrastructures.
@@ 2 Data Generator DCSim
It has been shown that after calibration, DCSim can achieve valid descriptions of real-world computing workflow executions and systems Horzela2023Proceeding ; Horzela2023_1000165566 . However, the execution time of DCSim scales superlinearly with the size of the simulated computing system, which limits its feasibility and necessitates workarounds in studies concerning computing systems at a global scale, like the WLCG. We therefore utilize DCSim as a baseline for testing ML surrogates in simulation.
@@ 3.2 Model Training
We split the data into training and evaluation datasets to train the model, using a 70:30 ratio per simulation of the same length. We employ sequence-to-sequence prediction, predicting exactly one output per input row. The mean squared error (MSE) loss function is used to penalize larger errors more heavily than smaller ones, emphasizing the importance of minimizing significant discrepancies between predicted and actual values.
@@ 4.1 Homogeneous Jobs
All models exhibit a negative R-squared value for compute_time, indicating a worse fit than a naive baseline. The KDE plots show that the target distribution exhibits two prominent spikes for the compute_time. While the BiGRU and BiLSTM models successfully capture the general structure of the distribution, they fail to emphasize these two spikes sufficiently. The transformer model, however, predicts mainly the mean values for compute time, failing to capture the actual distribution. While the R-squared value for input_files_transfer_time is positive, it is close to zero for all models, indicating only a slightly better fit than the naive baseline. The KDE plots show that the models capture some distribution features again but fail to emphasize the actual shape. In summary, all models failed to predict the observables beyond the first statistical moment, potentially due to the lack of input features providing additional context constraining the job execution, particularly information about the platform architecture. Although the size of the modelled infrastructure is rather minimal in this example, the surrogate ansatz already leads to improvements in the execution time of minimum two orders of magnitude on our test systems ($\mathcal{O}(10\,\mathrm{s})\to\mathcal{O}(100\,\mathrm{ms})$).